If a 19.5 mL sample of 1.5 M solution of each of the following acids is reacted with 0.80 M NaOH, how many milliliters of the NaOH are required for the titration? What is the total volume (in mL) of solution at the equivalence point?


19.5 mL of H2SO4 titrated with 0.80 M NaOH:

volume of NaOH ______ mL

total volume______ mL

19.5 mL of HC2H3O2 titrated with 0.80 M NaOH:

volume of NaOH_____ mL

total volume ____mL

19.5 mL of HCl titrated with 0.80 M NaOH:

volume of NaOH_______ mL

total volume ______mL
Answer #1

H2SO4(aq) + 2NaOH(aq) --------------> Na2SO4(aq) + 2H2O(l)

1 mole           2moles

H2So4                                              NaOH

M1   = 1.5M                                     M2   = 0.80M

V1   = 19.5ml                                    V2   =

n1    = 1                                           n2 = 2

            M1V1/n1   = M2V2/n2

                 V2      = M1V1n2/M2n1

                             = 1.5*19.5*2/(0.80*1)

                               = 73.125ml

volume of NaOH   = 73.125ml

total volume = 73.125 + 19.5   = 92.625ml

CH3COOH(aq) + NaOH(aq) --------------> CH3COONa(aq) + H2O(l)

1 mole           1moles

CH3COOH                                              NaOH

M1   = 1.5M                                     M2   = 0.80M

V1   = 19.5ml                                    V2   =

n1    = 1                                           n2 = 1

            M1V1/n1   = M2V2/n2

                 V2      = M1V1n2/M2n1

                             = 1.5*19.5*1/(0.80*1)

                             = 36.5625ml

volume of NaOH   = 36.5625ml

total volume = 36.5625 + 19.5   = 56.0625ml

HCl(aq) + NaOH(aq) -------------->NaCl(aq) + H2O(l)

1 mole           1moles

HCl                                                NaOH

M1   = 1.5M                                     M2   = 0.80M

V1   = 19.5ml                                    V2   =

n1    = 1                                           n2 = 1

            M1V1/n1   = M2V2/n2

                 V2      = M1V1n2/M2n1

                             = 1.5*19.5*1/(0.80*1)

                             = 36.5625ml

volume of NaOH   = 36.5625ml

total volume = 36.5625 + 19.5   = 56.0625ml